OpenStudy (anonymous):

Evaluate k7 ÷ k3, when k = 8

OpenStudy (anonymous):

(8)7÷(8)3= 56 ÷ 24= 2.333...

OpenStudy (anonymous):

k7=56 k3=24 So it would be 56/24= 2.33